b) Find the solution of the Riccati equation
dy/dx=(2cos^2x-sin^2x+y^2)/2cosx ;y(x)=sinx
1
Expert's answer
2020-03-24T11:41:39-0400

The general for Ricatti's equation:

"y'=P(x)+Q(x)y+R(x)y^2"

Here

"P(x)=\frac{2\cos^2x-\sin^2x }{2\cos x}; Q(x)=0;R(x)=\frac{1}{2\cos x}"

Since "y_1=\sin x" is the particular solution of the equation, we can build the general solution of the equation in form

"y=y_1+\frac{1}{w(x)}"

Where "w(x)" is the solution to the first-order linear equation

"w'=-(Q(x)+2R(x)y_1)w-R(x)"

or

"w'=-\frac{\sin x}{\cos x}w-\frac{1}{2\cos x}"

"w'+w\tan x=-\frac{1}{2\cos x}"

This is a first order linear equation, which solution is

"w=c_1\cos x-\frac{\sin x}{2}"

Therefore, the general solution of the equation is

"y=\sin x+\frac{2}{C\cos x-sinx}, C=2c_1"